# Probing the many-body localized spin-glass phase through quench dynamics

### v1

@@ Abstract
More recently, increasing attention has been devoted to the interplay of disorder and topological order [27, 28, 29, 30, 31, 32, 33, 34, 35], as shown e.g. in the interacting Majorana chain. In these systems, the disorder of the local potential and of the bond coupling compete, and two distinct MBL phases emerge [36, 27, 28, 29, 30]. When the on-site disorder dominates, the system is in a paramagnetic MBL phase [30, 28, 29], where the original topological order is completely destroyed and the system behaves according to standard MBL: area-law entanglement of eigenstates [37, 38], Poissonian level statistics [39], and slow entanglement growth [40]. However, in the opposite scenario of dominating bond disorder the system enters a novel MBL phase, characterized by long-range correlations, degeneracies in the spectrum and a cat-like structure of the eigenstates [27, 30, 28, 29] leading to higher entanglement as compared to the paramagnet MBL. In spite of the growing interest, the dynamical characterization of this MBL spin-glass (MBL-SG) phase is still limited to small systems [32].
@@
In the remainder of this work we will investigate the dynamics deep in the spin-glass phase, where we can provide an analytical theory explaining our numerical observations. We will focus on the dynamics from random product states in the computational basis that can be easily prepared and probed experimentally in a variety of quantum simulation platforms [13, 42].
@@
Correlation length.— To additionally characterize the fate of the MBL spin-glass regime, we investigate the behavior of the spin-spin correlation function averaged over disorder realizations, $\overline{\langle S^{x}_{i}S^{x}_{j}\rangle}=\overline{\langle\psi(t)|S^{x}_{i}S^{x}_{j}|\psi(t)\rangle}$. We observe that the averaged correlation function decays exponentially with distance, $\overline{\langle S^{x}_{i}S^{x}_{j}\rangle}\propto\exp(-|i-j|/\xi(t))$, defining a time-dependent correlation length $\xi(t)$. By fitting the correlation function of the central spin, $i=N/2$, to an exponential decay with distance $j$, we extract the time-dependent correlation length $\xi(t)$ and examine its evolution for different values of $\delta$ (see [45] for fit examples).
@@
More broadly, the slow growth of entanglement we observe is qualitatively consistent with the dynamical RG treatment constructed for a different spin-1/2 model and a specific Neél initial state [43]. Although our analytical results can serve as a starting point for extending the dynamical RG of Ref. [43] to our model, this remains an open question for future work. Our work shows that phenomenological signatures of an infinite randomness fixed point can be detected in quench dynamics. Extending these to the direct observation of clusters of decimated spins as inert degrees of freedom entangled among themselves but decoupled from the rest of the system, and of their hierarchical structure is a fascinating direction for further work. Finally, our results show that the MBL-SG phase can also be dynamically probed in quantum simulation platforms, since non-trivial dynamics arise at sufficiently early times.
@@ Acknowledgements.
Acknowledgments – We thank D. A. Abanin for insightful discussions in the early stages of this work. P. B. acknowledges support by the Austrian Science Fund (FWF) (Grant Agreement No. 10.55776/ESP9057324). The authors acknowledge support by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 850899). M. L. acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2111 – 390814868. The Authors acknowledge PRACE for awarding access to Joliot-Curie at GENCI@CEA, France, where the TEBD simulations were performed. The TEBD simulations were performed using the ITensor library [50].
@@ Appendix C Examples of correlation function exponential fit and power law fit for $\xi(t)$
In the main text we defined the corelation length as the length scale for the exponential decay of the disorder averaged correlation function $\overline{\langle S^{x}_{i}S^{x}_{j}\rangle}\propto e^{-|i-j|/\xi}$. We notice that for our particular choice of initial state $\langle\psi(t)|S^{x}_{j}|\psi(t)\rangle=0,\;\forall j,t$ and therefore the connected part of the correlation function coincides with $\langle S^{x}_{i}S^{x}_{j}\rangle$. In an infinite system, the disorder averaged correlation length is expected to be the same on all sites, as each local disorder is independently drawn. However, as we deal with finite systems, we chose to study the decay of the correlation function from the central spin $i=N/2$ to reduce the finite size effects. This defines the correlation length $\xi_{N/2}$ shown in the main text, which upon sufficient disorder averaging coincides with the generic correlation length $\xi$.

### v2

@@ Abstract
More recently, increasing attention has been devoted to the interplay of disorder and topological order [27, 28, 29, 30, 31, 32, 33, 34, 35, 36], as shown e.g. in the interacting Majorana chain. In these systems, the disorder of the local potential and of the bond coupling compete, and two distinct MBL phases emerge [37, 27, 28, 29, 30]. When the on-site disorder dominates, the system is in a paramagnetic MBL phase [30, 28, 29], where the original topological order is completely destroyed and the system behaves according to standard MBL: area-law entanglement of eigenstates [38, 39], Poissonian level statistics [40], and slow entanglement growth [41]. However, in the opposite scenario of dominating bond disorder the system enters a novel MBL phase, characterized by long-range correlations, degeneracies in the spectrum and a cat-like structure of the eigenstates [27, 30, 28, 29] leading to higher entanglement as compared to the paramagnet MBL. In spite of the growing interest, the dynamical characterization of this MBL spin-glass (MBL-SG) phase is still limited to small systems [32].
@@
In the remainder of this work we will investigate the dynamics deep in the spin-glass phase, where we can provide an analytical theory explaining our numerical observations. We will focus on the dynamics from random product states in the computational basis that can be easily prepared and probed experimentally in a variety of quantum simulation platforms [13, 43]. Unless otherwise specified, the data reported are obtained averaging over $50$ random initial states, each corresponding to a different disorder realization.
@@
Correlation length.— To additionally characterize the fate of the MBL spin-glass regime, we investigate the behavior pf the absolute value of the connected spin-spin correlation function averaged over disorder realizations and initial states, $\overline{|\langle S^{x}_{i}S^{x}_{j}\rangle|}=\overline{|\langle\psi(t)|S^{x}_{i}S^{x}_{j}|\psi(t)\rangle-\langle\psi(t)|S^{x}_{i}|\psi(t)\rangle\langle\psi(t)|S^{x}_{j}|\psi(t)\rangle|}$. We observe that the averaged correlation function decays exponentially with distance, $\overline{|\langle S^{x}_{i}S^{x}_{j}\rangle|}\propto\exp(-|i-j|/\xi(t))$, defining a time-dependent correlation length $\xi(t)$. By fitting the correlation function of the central spin, $i=N/2$, to an exponential decay with distance $j$, we extract the time-dependent correlation length $\xi(t)$ and examine its evolution for different values of $\delta$ (see [46] for fit examples).
@@
More broadly, the slow growth of entanglement we observe is qualitatively consistent with the dynamical RG treatment constructed for a different spin-1/2 model and a specific Neél initial state [44]. Although our analytical results can serve as a starting point for extending the dynamical RG of Ref. [44] to our model, this remains an open question for future work. Our work shows that phenomenological signatures of an infinite randomness fixed point can be detected in quench dynamics. Extending these to the direct observation of clusters of decimated spins as inert degrees of freedom entangled among themselves but decoupled from the rest of the system, and of their hierarchical structure is a fascinating direction for further work [33]. Finally, our results show that the MBL-SG phase can also be dynamically probed in quantum simulation platforms, since non-trivial dynamics arise at sufficiently early times.
@@ Acknowledgements.
Acknowledgments – We thank D. A. Abanin for insightful discussions in the early stages of this work. P. B. acknowledges support by the Austrian Science Fund (FWF) [Grant Agreement No. 10.55776/ESP9057324]. This research was funded in whole or in part by the Austrian Science Fund (FWF) [10.55776/COE1]. The authors acknowledge support by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 850899). M. L. acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2111 – 390814868. The Authors acknowledge PRACE for awarding access to Joliot-Curie at GENCI@CEA, France, where the TEBD simulations were performed. The TEBD simulations were performed using the ITensor library [51].
@@ Appendix C Examples of correlation function exponential fit and power law fit for $\xi(t)$
In the main text we defined the corelation length as the length scale for the exponential decay of the disorder averaged absolute value of the correlation function $\overline{|\langle S^{x}_{i}S^{x}_{j}\rangle|}\propto e^{-|i-j|/\xi}$. We notice that for our particular choice of initial state $\langle\psi(t)|S^{x}_{j}|\psi(t)\rangle=0,\;\forall j,t$ and therefore the connected part of the correlation function coincides with $\langle S^{x}_{i}S^{x}_{j}\rangle$. In an infinite system, the disorder averaged correlation length is expected to be the same on all sites, as each local disorder is independently drawn. However, as we deal with finite systems, we chose to study the decay of the correlation function from the central spin $i=N/2$ to reduce the finite size effects. This defines the correlation length $\xi_{N/2}$ shown in the main text, which upon sufficient disorder averaging coincides with the generic correlation length $\xi$.